# AffordanceSAM: Segment Anything Once More in Affordance Grounding

### v1

@@ 4.2 Main Results
Figure 4: Qualitative results with novel objects from the internet. learned and novel affordance functions are marked in blue and yellow. Cross-View-AG and LOCATE can not generalize to new affordance functions, so we do not highlight any region. OOAL often outputs suboptimal affordance maps when encountering new objects and affordance functions. By contrast, our AffordanceSAM can still give reasonable affordance maps.
Qualitative Results on Internet Images. To further demonstrate the generalization capacity of AffordanceSAM, we also present qualitative comparisons when facing new objects selected from the internet. As we can see, AffordanceSAM not only produces very reasonable affordance maps for learned affordance functions showed in the first three rows of Figure 4, but is also able to handle novel actions presented in the last three rows of Figure 4. By contrast, understanding new actions is beyond the capability of Cross-View-AG and LOCATE, and OOAL can not make concentrated predictions for things beyond training data. These results substantiate that our AffordanceSAM can benefit so greatly from the foundation model SAM, thus showcasing leading affordance grounding performance towards novel objects and affordance functions.
Stage1 Stage2 Stage3 KLD $\downarrow$ SIM $\uparrow$ NSS $\uparrow$
@@
Table 3: Ablation results of coarse-to-fine training recipe. First, we gradually add the training data divided by stage. Then, we provide the results when directly combine all the data and train for one stage.
@@
Table 5: Ablation results of number of filtration. Num. indicates the number of times we filter the affordance maps shown in steps 4$\sim$6 of Algorithm 1.
@@
[…]
@@
4.3 Ablation Study
In this section, we conduct an extensive ablation study to reveal the contribution and design-space of each component. Unless otherwise specified, we report the results on hard split of AffordanceSAMSAM1 using our default settings. Please refer to Appendix 6 to see the detailed illustration of the evaluation setup of each table.
Contribution of Coarse-to-Fine Training Recipe. As depicted in Table 7, our proposed coarse-to-fine training recipe demonstrates a clear performance enhancement. It is noteworthy that directly employing EVF-SAM to assess results can be highly unsatisfactory (see the first row), which is consistent with the analysis in Section 1 and Figure 1 that SAM lacks the capability to recognize affordance verbs and thereby overly segments objects. Encouragingly, after our three consecutive stages of training in a stepwise manner, it can be found that the affordance performance has been progressively enhanced. Moreover, as shown in the last row, directly combining all the data can lead to worse result, this may because directly mixing different quality of data and ground-truth format causes the performance degeneration, which is consistent with the observations in tuning LLMs to MLLMs  [28, 9, 4].
Design Choices of Affordance-Adaption Module. We start from EVF-SAM without any structural changes and gradually integrate our methods to analyze the effect of each proposed design in affordance-adaption module. The results in Tab 7 reveal that each element consistently delivers notable improvements. In particular, we notice that adding learnable queries benefits a lot, which is in line with findings in other works [22, 57, 18] that use prompt learning when adapting a foundation model to downstream tasks.
Effect of Post-Processing Program. As we display in the first two columns of Figure 5, the quality of affordance maps directly output by LOCATE is sometimes terrible, which would result in corruption of the model capability supervised on these maps (see Table 7 when $\gamma=0$). By contrast, after processing with an optimal filtration intensity of $\gamma=0.45$ and number of 3 (see Table 7), most of the irrelevant areas are eliminated, while areas related to the functional part are preserved (see the second two columns of Figure 5). This optional design choose also leads to a significant performance improvement across all metrics.
5 Conclusion
In this paper, we introduce AffordanceSAM, a new state-of-the-art affordance grounding model. Built upon EVF-SAM with our proposed affordance-adaption module, AffordanceSAM demonstrates superior performance in affordance grounding tasks after a systematically designed coarse-to-fine training. The zero-shot affordance grounding evaluation on internet images further demonstrates the generalization capacity of AffordanceSAM. Despite AffordanceSAM still has its constraints, such as the suboptimal performance in complex scenarios (see Appendix 7), we still believe AffordanceSAM can become the new baseline for subsequent studies and offer timely insights into how to leverage and extend SAM-like foundation models to benefit more relevant vision tasks with limited labeled data resources.
@@ Appendix 1 More Discussion on Related Work
Affordance learning for robotics. Recently, Numerous methodologies have been developed by researchers to extract and interpret affordance information to enable the robots to grasp in a complex, dynamic environments [2, 49]. Specifically, several studies [1, 21, 30] leverage affordance to establish correlations between objects, tasks, and manipulation strategies for robotic grasping. Other studies [5, 6, 3] focus on deriving affordance knowledge from resources that can be deployed on real robotic systems. In this study, we only focus on the visual affordance grounding task, and our main insight is taking advantage of SAM’s generalization ability to compensate for the deficiencies of the previous methods. But we also believe our propose AffordanceSAM can greatly improve the robot’s ability to recognize and grasp objects and becomes a default option for robot deployment by other researchers.
Segment anything model and its extension. The Segment Anything Model (SAM) [20, 41] is an interactive segmentation framework that produces binary masks in response to various prompt types (points, boxes, and coarse masks). Trained on a comprehensive dataset, SAM exhibits robust generalization capacity for segmenting diverse objects. Building upon SAM, SAM-HQ [17] addresses the segmentation quality of SAM by introducing a set of hq-tokens; LISA [23] combines SAM with LLaVA [28] to enable segmentation model with reasoning ability. Our work also start from SAM, but we focus on converting SAM to be a generalized affordance grounding model with our proposed affordance-adaption module and coarse-to-fine training recipe.
@@ Appendix 2 More Details about Our Datasets
Table 8: Statistics of datasets we used in each stage. Part: part-level annotation. Obj.: number of object classes. Aff: number of affordance function classes. Img.: number of images.
Dataset Year Part Obj. Aff. Img.
Stage1
PADv2 [51] 2021 $\times$ 103 39 30,000
RGB-D Part Affordance [33] 2022 $\checkmark$ 37 15 47,210
Handal [13] 2023 $\checkmark$ 212 17 30,800
Stage2 and Stage3
AGD20k [29] 2021 $\checkmark$ 50 36 23,816
Here we provide the detailed dataset setup for training and testing AffordanceSAM. The statistics of each dataset is provided in Table 8.
Stage1. In this stage, we use PADv2 [51], Handal [13], and RGB-D Part Affordance [33]. All three datasets are composed of different object and affordance categories, and images of the same affordance and category are treated as one class. The ground-truth of three datasets is binary mask, but a slight different is that Handal and RGB-D Part Affordance label the part of the object according to the affordance action, and PADv2 only gives the mask of the whole obejct. Moreover, as Handal and RGB-D Part Affordance are obtained from continuous frames of videos, which means many of the images of same object category are very similar. We randomly sampled 5 images for those belong to a video clip to avoid potential over-fitting problem. And for PADv2, we filter duplicate images and add remaining images in our training data. Noticing that we do not split any object in this stage like in stage2 and stage3, because the output form of this stage is completely different from the output used in the final evaluation.
Stage2 and stage3. In these two stages, we strictly follow AffordanceLLM [39] to split AGD20K [29] for training and testing our model. The easy split is the original unseen split of AGD20K, which has a lot of similarities between the objects in the train and test sets. The hard split ensures that there is no overlap between the object categories in the train and test set. This setting can help to reflect the model generalization ability when the similarity between the training data and the test data changes. More details can be found in AffordanceLLM’s paper and project [^2] https://jasonqsy.github.io/AffordanceLLM/.
@@
Total batch size $32$ $32$ $32$
@@
Encoder&\&&Decoder SAM-ViT-H [20] SAM2-hiera-L [41]
Image input size 1024$\times$1024 1024$\times$1024
Patch size 16 16
Encoder hidden dimension 1280 576
Global attention blocks [7, 15, 23, 31] [23, 33, 43]
@@
Encoder BEIT-3-L [48] BEIT-3-L [48]
Text tokenizer XLM-Roberta [10] XLM-Roberta [10]
Image input size 224$\times$224 224$\times$224
Patch size 16 16
Encoder hidden dimension 1024 1024
@@ Appendix 3 More Details about Implementation
In this section, we provide specific training hyperparameters in each stage and model configuration of different version of AffordanceSAM in Table 9 and Table 10, respectively.
What’s more, we use different data augmentation when training AffordanceSAMSAM1 (large-scale jitter [11] with a scale range of [0.1, 2.0], rand-hflip with a ratio of 0.5) and AffordanceSAMSAM2 (only rand-hflip with a ratio of 0.5), because we find that using the data augmentation in accordance with that used in the backbone’s pre-training [20, 41] can lead to better results. We use the checkpoint after the whole training in first two stages as the initialization of the next stage.
@@ Appendix 4 Details of Each Metrics for Evaluation
In this section, we explain the metrics (KLD [7], SIM [46], and NSS [37]) to evaluate models.
@@
$\mathrm{SIM}\left(M,M^{\prime}\right)=\sum_{i}\min\left(M_{i},M^{\prime}_{i}\right),\$ (7)
@@ Appendix 5 Methods for Comparison
Affordance grounding methods can be mainly summarized into two main categories: weakly supervised and fully supervised methods. For weakly supervised models, they do not train on explicit labels of the affordance map. Instead, they are trained on two views (egocentric and exocentric) of the same object. As for fully supervised models, they are supervised under dense labels. In what follows, we explain the main idea of baseline methods that we used for the evaluation and comparison.
Cross-View-AG [29] proposes a novel framework to extract invariant affordance from exocentric interactions and transfer it to egocentric views. This includes an Affordance Invariance Mining module to minimize intra-class differences in exocentric images and an Affordance Co-relation Preserving strategy to align co-relation matrices between views for affordance perception and localization.
LOCATE [24] first localizes where the interaction happens and identifies interaction regions in exocentric views, then uses a PartSelect module to extract affordance-specific information, and transfers this knowledge to egocentric views for affordance grounding using only image-level labels.
AffordanceLLM [39] leverages LLaVA [28] with a new mask token similar to LISA [23] and introduce depth maps as 3D information to build the main pipeline. A light-weight mask decoder trained from scratch is also introduced to produce affordance map.
OOAL [25] proposes a vision-language framework that includes CLIP [40] and DINOv2 [35]. To boost the alignment between visual features of DINOv2 and affordance text embeddings extracted by CLIP, several light-weight modules like learnable text prompt tokens and a CLS-guided transformer decoder are introduced.

### v2

@@ 3 Labeled Part of AGD20K Human Annotated Affordance Map 1135 (Easy Split)
Figure 5: Qualitative results with novel objects from the internet. learned and novel affordance actions are marked in blue and yellow. Cross-View-AG and LOCATE can not generalize to new affordance actions, so we do not highlight any region. OOAL often outputs suboptimal affordance maps when encountering new objects and affordance actions. By contrast, our AffordanceSAM can still give reasonable affordance maps.
Experiments
Experimental Setup
Implementation Details. In stage1, we initialize our AffordanceSAM with EVF-SAM (Zhang et al. 2024) and train for 13 epochs with base learning as 0.00002. In stage2, we use the weights obtained from the first stage and randomly initialize affordance-adaption module, we train the model with the same number of epochs and base learning as stage1. In stage3, we further train our model for 26 epochs with larger base learning as 0.00004. In all training stages, we employ an AdamW optimizer (Kingma 2014) with a cosine learning rate scheduler, and the total batchsize per iteration is 32 using 4 NVIDIA A100 (80GB) GPUs.
Evaluation. We strictly follow the settings in AffordanceLLM to have two splits (easy split and hard split) of AGD20K to test our model. The easy split is the original unseen split of AGD20K, which has a lot of similarities between the objects in the train and test sets. The hard split ensures that there is no overlap between the object categories in the train and test set (both in our stage 2 and stage 3). Meanwhile, we follow the metrics provided in AGD20K to evaluate our model, which is Kullback-Leibler Divergence (KLD) (Bylinskii et al. 2018), Similarity (SIM) (Swain and Ballard 1991) and Normalized Scanpath Saliency (NSS) (Peters et al. 2005).
Methods for Comparison. Affordance grounding methods can be mainly summarized into two main categories: weakly supervised and fully supervised methods. We compare our approach against representative weakly supervised baselines (Cross-View-AG (Luo et al. 2022), LOCATE (Li et al. 2023a), R-Mamba (Wang et al. 2025), and PLSP (Xu and Mu 2025)), and fully supervised baselines (AffordanceLLM (Qian et al. 2024) and OOAL (Li et al. 2024)).
[…]
Main Results
We present both quantitative and qualitative experimental results and analyses of our proposed method. Below, we summarize the key findings:
Effectiveness of AffordanceSAM Model and C2F-Aff Dataset. As illustrated in Table 2, the AffordanceSAM model demonstrates remarkable performance, achieving competitive results even without leveraging a larger dataset (C2F-Aff). This is particularly notable when compared to weakly supervised methods. Trained only the AGD20K-Fully, AffordanceSAM outperforms AffordanceLLM and outperforms the previous SOTA OOAL model in terms of KLD and NSS, across both the easy and hard splits. From a data perspective, the introduction of the C2F-Aff dataset has significantly enhanced the performance of all fully supervised methods in the challenging hard splits, and the improvement is most prominent for our model. It is worth noting that all the fully supervised methods trained on C2F-Aff dataset outperformed other weakly supervised candidates. This illustrates the potential of the fully supervised method when scaling the supervised data.
Generalized Performance when Combining the Model and Dataset. By combining the model with the rich annotations of C2F-Aff, AffordanceSAM achieves state-of-the-art (SOTA) performance across most metrics. Specifically, AffordanceSAM attains the highest SIM and NSS scores on the easy and hard splits, respectively. Furthermore, AffordanceSAM demonstrates exceptional generalization to novel objects and actions, as evidenced by qualitative results on both AGD20K images and internet images (Figure 4 and Figure 5). For instance, when handling the object “cup”, AffordanceSAM demonstrates a superior understanding of affordance actions compared to other methods and generates more accurate and precise affordance maps according to different actions “hold” and “drink”. These results substantiate that our AffordanceSAM can benefit so greatly from the foundation model SAM, as well as the proposed C2F-Aff dataset, thus showcasing leading affordance grounding performance towards novel objects and affordance actions.
Ablation Study
In this section, we conduct an extensive ablation study to reveal the contribution and design-space of each component. Unless otherwise specified, we report the results on hard split of AffordanceSAM using our default settings. Please refer to Appendix 6 to see the detailed illustration of the evaluation setup of each table.
Part 1 Part 2 Part 3 KLD $\downarrow$ SIM $\uparrow$ NSS $\uparrow$
@@
Table 3: Ablation results of coarse-to-fine dataset and training recipe. First, we gradually add the training data of each part in C2F-Aff . Then, we provide the results when directly combine all the data and train for one stage.
@@
[…]
@@
w/ $LQ,F_{t},F_{v}\text{ w/ m}$ 1.128 0.514 1.761
w/ $LQ,F_{t},F_{v}\text{ wo/ m}$ 1.192 0.503 1.749
Table 4: Ablation results of affordance-adaption module. LQLQ: learnable queries. $F_{t}$: interaction with text feature. $F_{v}\text{ w/ m}$: interaction with fused image features. $F_{v}\text{ w/o m}$: interaction with only last layer image feature.
@@
Contribution of Coarse-to-Fine Dataset and Training Recipe. As depicted in Table 7, our proposed coarse-to-fine dataset (C2F-Aff) and training recipe demonstrate a clear performance enhancement. It is noteworthy that directly employing EVF-SAM to assess results can be highly unsatisfactory (see the first row), which is consistent with the analysis in Introduction and Figure 2 that most vision foundation models like SAM lack the capability to recognize affordance verbs and thereby overly segments objects. Encouragingly, after trained on our three consecutive part of C2F-Aff in a stepwise manner, it can be found that the affordance performance has been progressively enhanced. Moreover, as shown in the last row, directly combining all the data can lead to worse result, this may because directly mixing different quality of data and ground-truth format causes the performance degeneration, which is consistent with the observations in tuning LLMs to MLLMs  (Liu et al. 2023; Chen et al. 2024; Bai et al. 2023).
@@
Design Choices of Affordance-Adaption Module. We start from EVF-SAM without any structural changes and gradually integrate our methods to analyze the effect of each proposed design in affordance-adaption module. The results in Tab 7 reveal that each element consistently delivers notable improvements. In particular, we notice that adding learnable queries benefits a lot, which is in line with findings in other works (Lafon et al. 2024; Zhou et al. 2022; Khattak et al. 2023) that use prompt learning when adapting a foundation model to downstream tasks.
Effect of Post-Processing Program. As we display in the first two columns of Figure 6, the quality of affordance maps directly output by LOCATE is sometimes terrible, which would result in corruption of the model capability supervised on these maps (see Table 7 when $\gamma=0$). By contrast, after processing with an optimal filtration intensity of $\gamma=0.45$ and a number of 3 (see Table 7), most of the irrelevant areas are eliminated, while areas related to the functional part are preserved (see the second two columns of Figure 6). This optional design choice also leads to a significant performance improvement across all metrics.
Conclusion
In this paper, we introduce AffordanceSAM, a new state-of-the-art (SOTA) affordance grounding model. Built upon EVF-SAM with our proposed affordance-adaption module and trained with our systematically curated C2F-Aff dataset in a coarse-to-fine training manner, AffordanceSAM demonstrates superior performance in affordance grounding tasks. The zero-shot affordance grounding evaluation on internet images further demonstrates the generalization capacity of AffordanceSAM. We believe AffordanceSAM can become the new baseline, and the C2F-Aff dataset can become an effective data resource for subsequent studies and offer timely insights into how to leverage and extend SAM-like foundation models to benefit more relevant vision tasks.
@@ Appendix 1 More Discussion on Related Work
Affordance Learning for Robotics. Recently, Numerous methodologies have been developed by researchers to extract and interpret affordance information to enable the robots to grasp in a complex, dynamic environments (Ardón et al. 2020; Yang et al. 2023). Specifically, several studies (Ardón et al. 2019; Kokic et al. 2017; Ma et al. 2024) leverage affordance to establish correlations between objects, tasks, and manipulation strategies for robotic grasping. Other studies (Bharadhwaj, Gupta, and Tulsiani 2023; Borja-Diaz et al. 2022; Bahl et al. 2023) focus on deriving affordance knowledge from resources that can be deployed on real robotic systems. In this study, we only focus on the visual affordance grounding task, and our main insight is taking advantage of SAM’s generalization ability to compensate for the deficiencies of the previous methods. But we also believe our proposed AffordanceSAM can greatly improve the robot’s ability to recognize and grasp objects and become a default option for robot deployment by other researchers.
Segment anything model and its extension. The Segment Anything Model (SAM) (Kirillov et al. 2023; Ravi et al. 2024) is an interactive segmentation framework that produces binary masks in response to various prompt types (points, boxes, and coarse masks). Trained on a comprehensive dataset, SAM exhibits robust generalization capacity for segmenting diverse objects. Building upon SAM, SAM-HQ (Ke et al. 2024) addresses the segmentation quality of SAM by introducing a set of hq-tokens; LISA (Lai et al. 2024) combines SAM with LLaVA (Liu et al. 2023) to enable a segmentation model with reasoning ability. Our work also starts from SAM, but we focus on converting SAM to be a generalized affordance grounding model with our proposed affordance-adaption module and C2F-Aff dataset and training recipe.
@@ Appendix 2 More Details about Our Datasets
Here we provide the detailed dataset setup for C2F-Aff training and AGD20K testing of AffordanceSAM. The statistics of each dataset are provided in Table 8. Just as shown in this table and Table 1 of the main paper, our C2F-Aff dataset is divided into three parts according to the label format, and we will introduce each part in sequence parts.
Table 8: Statistics of datasets we used in for our C2F-Aff data. PL.: part-level annotation. Obj.: number of object classes. Aff: number of affordance action classes. Img.: number of images.
Dataset Year PL Obj. Aff. Img.
Part 1
PADv2 (Zhai et al. 2022) 2021 $\times$ 103 39 30,000
RGB-D Part Affordance (Myers et al. 2015) 2022 $\checkmark$ 37 15 47,210
Handal (Guo et al. 2023) 2023 $\checkmark$ 212 17 30,800
Part 2 and Part 3
AGD20k (Luo et al. 2022) 2021 $\checkmark$ 50 36 23,816
@@
Total batch size 3232 3232 3232
@@
Encoder&\&Decoder SAM-ViT-H (Kirillov et al. 2023)
Image input size $\times1024$
Patch size 16
Encoder hidden dimension 1280
Global attention blocks [7, 15, 23, 31]
@@
Encoder BEIT-3-L (Wang et al. 2023b)
Text tokenizer XLM-Roberta (Conneau 2019)
Image input size $\times224$
Patch size 16
Encoder hidden dimension 1024
@@
Part 1. In this part, we use PADv2 (Zhai et al. 2022), Handal (Guo et al. 2023), and RGB-D Part Affordance (Myers et al. 2015). All three datasets are composed of different object and affordance categories, and images of the same affordance and category are treated as one class. The ground-truth of three datasets is binary mask, but a slight different is that Handal and RGB-D Part Affordance label the part of the object according to the affordance action, and PADv2 only gives the mask of the whole object. Moreover, as Handal and RGB-D Part Affordance are obtained from continuous frames of videos, which means many of the images of the same object category are very similar. We randomly sampled 5 images for those that belong to a video clip to avoid a potential over-fitting problem. And for PADv2, we filter duplicate images and add the remaining images to our training data. Noticing that we do not split any object when training, because the output form of this stage is completely different from the output used in the final evaluation.
Part 2 and Part 3. In these two data parts, we use the labeled and unlabeled parts of AGD20K (Luo et al. 2022) and use LOCATE (Li et al. 2023a) to annotate the unlabeled samples. We have already discussed this point in detail in the main paper. Moreover, in order to test the model’s performance accurately and fairly, we strictly follow AffordanceLLM (Qian et al. 2024) to split AGD20K for training and testing our model. The easy split is the original unseen split of AGD20K, which has a lot of similarities between the objects in the train and test sets. The hard split ensures that there is no overlap between the object categories in the training and testing sets. This setting can help to reflect the model’s generalization ability when the similarity between the training data and the test data changes. More details can be found in AffordanceLLM’s paper and project.
@@ Appendix 3 More Details about Implementation
In this section, we provide specific training hyperparameters in each stage and model configuration of AffordanceSAM in Table 9 and Table 10, respectively. We use the checkpoint after the whole training in first two stages as the initialization of the next stage.
@@ Appendix 4 Details of Each Metrics for Evaluation
In this section, we explain the metrics (KLD (Bylinskii et al. 2018), SIM (Swain and Ballard 1991), and NSS (Peters et al. 2005)) to evaluate models.
@@
$\mathrm{SIM}\left(M,M^{\prime}\right)=\sum_{i}\min\left(M_{i},M^{\prime}_{i}\right),\\$ (7)
@@ Appendix 5 Methods for Comparison
Affordance grounding methods can be mainly summarized into two main categories: weakly supervised and fully supervised methods. For weakly supervised models, they do not train on explicit labels of the affordance map. Instead, they are trained on two views (egocentric and exocentric) of the same object. As for fully supervised models, they are supervised under dense labels. In what follows, we explain the main idea of the baseline methods that we used for the evaluation and comparison.
Cross-View-AG (Luo et al. 2022) proposes a novel framework to extract invariant affordance from exocentric interactions and transfer it to egocentric views. This includes an Affordance Invariance Mining module to minimize intra-class differences in exocentric images and an Affordance Co-relation Preserving strategy to align correlation matrices between views for affordance perception and localization.
LOCATE (Li et al. 2023a) first localizes where the interaction happens and identifies interaction regions in exocentric views, then uses a PartSelect module to extract affordance-specific information, and transfers this knowledge to egocentric views for affordance grounding using only image-level labels.
R-Mamba (Wang et al. 2025) first extracts feature embeddings from exocentric and egocentric images to construct the hypergraphs. Then, a Hypergraph-guided State Space (HSS) block is introduced to reorganize these local relationships from a global perspective to locate the affordance regions.
PLSP (Xu and Mu 2025) introduces a label refining stage, a fine-grained feature alignment process, and a lightweight reasoning module to boost the performance of weakly supervised affordance learning.
AffordanceLLM (Qian et al. 2024) leverages LLaVA (Liu et al. 2023) with a new mask token similar to LISA (Lai et al. 2024) and introduce depth maps as 3D information to build the main pipeline. A light-weight mask decoder trained from scratch is also introduced to produce an affordance map.
OOAL (Li et al. 2024) proposes a vision-language framework that includes CLIP (Radford et al. 2021) and DINOv2 (Oquab et al. 2023). To boost the alignment between visual features of DINOv2 and affordance text embeddings extracted by CLIP, several light-weight modules like learnable text prompt tokens and a CLS-guided transformer decoder are introduced.